CS 791 - Computational Geometry

K. Subramani LDCSEE, West Virginia University, Morgantown, WV ksmani@csee.wvu.edu

1 Policies and Procedures

(i) All homeworks must be turned in on the prescribed date. (ii) Homeworks must be written up in LaTeX. (iii) A maximum of 5 bonus points is reserved for class performance. (iv) Each presentation should be self-contained and include the following information:

(a) Statement of Problem,

(b) Related Work,

(c) Current contribution,

(d) Open problems.

(v) If your presentation can be converted into publishable work, you will be awarded an A, regardless of your homework performance.

(vi) The standard Computational Geometry conferences are: (a) ACM Symposium on Computational Geometry,

(b) Canadian Conference on Computational Geometry,

(c) Japanese Conference on Discrete and Computational Geometry. (vii) The standard Computational Geometry journals are:

(a) Computational Geometry: Theory and Applications,

(b) Discrete and Computational Geometry,

(c) International Journal of Computational Geometry and Applications. (viii) [GO97] contains an excellent compilation of recent results in Computational Geometry.

References

[GO97] Jacob E. Goodman and Joseph O’Rourke, editors. Handbook of Discrete and Computational Geometry. CRC Press, 1997.